Solve the equation
x^2y^2(2ydx+xdy)-(5ydx+7xdy)=0
1
Expert's answer
2020-09-22T16:28:07-0400

The above equation can be written as

"(2y^3x^2dx+x^3y^2dy)-(5ydx+7xdy)=0"

Reforming the equation as

"(y^3x^2dx+x^3y^2dy+y^3x^2dx)-(5ydx+7xdy)=0"

Multiplying and dividing by 3

"\frac{1}{3}(3y^3x^2dx+3x^3y^2dy)+y^3x^2dx-(5ydx+7xdy)=0"

Let total differential be "x^3y^3"

the "d(x^3y^3)=3x^2y^3dx+3y^2x^3dy"


"\frac{1}{3}d(x^3y^3)+x^2y^3dx-(5ydx+7xdy)=0"


"\frac{1}{3}d(x^3y^3)+x^2y^3dx-(5ydx+5xdy+2xdy)=0"


"\frac{1}{3}d(x^3y^3)+x^2y^3dx-5(ydx+xdy)-2xdy=0"

Lets select the total differential be xy then,

"d(xy)=xdy+ydx"

"\frac{1}{3}d(x^3y^3)+x^2y^3dx-5d(xy)-2xdy=0"


Consider the equation

"x^2y^3dx-2xdy=0"

Above equation can be written as,

"\int xdx=\int\frac{2}{y^3}dy"

Integrate both th sides

"\frac{x^2}{2}=-\frac{1}{y^2}+c"


Where c i a constant

Hence the overall solution of given equation is,

"\frac{x^3y^3}{3}-5xy+\frac{x^2}{2}+\frac{1}{y^2}=c"